$SPAD/src/input e02akf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
np1:=7
xmin:=-0.5
xmax:=2.5
a:Matrix SF:=
 [[2.53213 ,1.13032 ,0.27150 ,0.04434 ,0.00547 ,0.00054 ,0.00004 ]]
ia1:=1
la:=7
x:=-0.5
result:=e02akf(np1,xmin,xmax,a,ia1,la,x, -1)

———-

References^{[1] nothing}

4